# Diffusion Transformer Captures Spatial-Temporal Dependencies: A Theory for Gaussian Process Data

### v1

@@ 3.3 Bounding Correlation Truncation Error $\mathcal{E}_{2}$
Let $\mathbf{e}_{1},\dots,\mathbf{e}_{N}\in\mathbb{R}^{d_{e}}$ be $d_{e}$-dimensional time embedding of $h_{1},\dots,h_{N}$ with $\left\lVert\mathbf{e}_{i}\right\rVert_{2}=r$ for each $i$ such that there exists a positive and strictly increasing function $f$ with $\left\lVert\mathbf{e}_{i}-\mathbf{e}_{j}\right\rVert_{2}=f(|i-j|)\geq c{\left|i-j\right|}$ for an absolute constant $c>0$. Further, the covariance function $\gamma$ satisfies
@@
Firstly, Assumption 1 says that the time embedding (a.k.a. position embedding) preserves the gap between real times. For example, a common time embedding used in sequential data modeling (Vaswani et al., 2017) is sinusoidal transformations, where $\mathbf{e}_{i}=[r\sin(2i\pi/C),r\cos(2i\pi/C)]^{\top}\in\mathbb{R}^{2}$ for a positive radius $r$ and a constant $C>0$. We can check that $\left\lVert\mathbf{e}_{i}-\mathbf{e}_{j}\right\rVert_{2}=2r\sin({\left|i-j\right|}\pi/C)\geq 4r{\left|i-j\right|}/C$ is positive and approximately linearly increasing for a sufficiently large $C$.
Secondly, the covariance function decays exponentially fast and the speed is controlled by the exponent $\nu$ and the bandwidth $\ell$. Large $\ell$ or small $\nu$ indicates that the correlation between different times decays relatively slowly. Thus, the sequential data has some long-horizon dependencies. The range of $\nu$ includes the well-known quadratic-exponential (Gaussian) covariance function ($\nu=2$). Moreover, when $\nu=1$, the covariance function coincides with the correlation in a Brownian motion. Varying $\nu\in[1,2]$ can capture abundant temporal dependency patterns.
@@ 4 Approximation Theory of Score Function using Transformers
The proof is deferred to Appendix B.1. Here we observe that the approximation error depends on $\sigma_{t}$, which matches existing results for studying score approximation using neural networks (Oko et al., 2023; Chen et al., 2023a; Tang and Yang, 2024). Yet we remark that our algorithm unrolling approach is very different from existing works. Our results also hold for softmax activated transformer architectures, which is discussed in Appendix B.5. We remark on other interpretations.
Figure 3: Construction of score function approximation using a transformer. By rewriting the score function as the optimizer of a quadratic objective function, we use gradient descent algorithm to approximate the optimizer. We allow correlation truncation to manipulate the maximum length of temporal dependencies to model in Lemma 1. Then we construct a transformer architecture to unroll the gradient descent algorithm for score approximation in Theorem 1. Each gradient descent iteration is realized by a multiplication module followed by two transformer blocks. In the first transformer block, its attention layer calculates correlation $\bm{\Gamma}_{ij}$ utilizing time embedding. The second transformer block calculates the linear offset $-\eta_{t}\sigma_{t}^{2}\mathbf{s}+\eta_{t}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})$ in (4).
@@ 5 Sample Complexity of Diffusion Transformer
However, this loss function is not directly implementable due to $\nabla\log p_{t}$ being unknown and numerical instability when $t$ approaches zero. Therefore, we consider the following loss for practical implementation,
@@
Suppose Assumption 1 holds with $\ell\leq c^{\nu}$. Assume there exists a constant $C>1$ such that $C^{-1}\leq\lambda_{\min}(\bm{\Sigma})\leq\lambda_{\max}(\bm{\Sigma})\leq C$ and $\left\lVert\bm{\mu}\right\rVert_{\infty}\leq C$. We choose the transformer architecture ${\mathcal{T}}(D,L,M,B,R_{t})$ as in Theorem 1 with $\epsilon=1/n$. By setting the terminal time $T=\log(n)$, considering $\widehat{P}$ generated by the empirical risk minimizer $\widehat{\mathbf{s}}$, we have
@@
The proof is deferred to Appendix C. We remark that the emergence of $\widetilde{P}_{0}$ owes to the early-stopping in score estimation, which is obtained by evolving $P_{0}$ along the forward process for time $t_{0}$. The optimal choice on $t_{0}$ depends on the condition number $\kappa_{t_{0}}$. For fast decay covariance functions, $\kappa_{t_{0}}=\mathcal{O}(1)$ and we can choose $t_{0}=1/n$ and the generalization error is in the order of $\widetilde{\mathcal{O}}(\sqrt{\ell Nd^{3}/n})$. In this case, we have a relatively weak dependence on $N$, demonstrating the efficiency of diffusion transformers in sequential data modeling. On the other hand, when $\kappa_{t_{0}}$ is large, i.e., with the presence of long-horizon dependencies, the learning efficiency suffers from heavier dependence on $N$. We experiment with various decaying patterns in Section 6 and show the corresponding performance.
@@ 6 Numerical Results
In this section, we conduct experiments on diffusion transformer for learning synthetic Gaussian process data. We test various aspects of our theory, including learning efficiency and the capture of spatial-temporal dependencies.
@@
We consider synthetic Gaussian process data with $d=8$ and $N=128$. We stick to the stationary setting, where the covariance matrix is generated by $\bm{\Sigma}=\mathbf{A}^{\top}\mathbf{A}\in\mathbb{R}^{8\times 8}$ for a Gaussian random matrix $\mathbf{A}\in\mathbb{R}^{8\times 8}$. We set the mean vector $\bm{\mu}=\mathbf{0}$ for simplicity. The covariance function is chosen as in Assumption 1 with $\gamma(h_{i},h_{j})=\exp\left(-|h_{i}-h_{j}|^{\nu}/\ell\right).$ We set $\ell$ and $\nu$ as hyperparameters and report their influences on the learning. We generate $n\in\{1000,3200,10000,32000,100000\}$ sequences from the Gaussian process as our training data in different settings. The diffusion transformer is implemented based on the DiT (Peebles and Xie, 2023) code base. We set the number of transformer blocks to be $12$ throughout all the experiments. We modify the patchify module in DiT to cope with our Gaussian process data. Additional implementation details can be found in Appendix D.
@@
We dive into transformer blocks to understand how attention layer captures dependencies. Inside a transformer block, the input is a concatenation of time embedding and a data vector written as $[\mathbf{z}_{i,t}^{\top},\mathbf{e}_{i}^{\top}]^{\top}\in\mathbb{R}^{32}$, where $\mathbf{e}_{i}\in\mathbb{R}^{16}$ is the time embedding and $\mathbf{z}_{i,t}\in\mathbb{R}^{16}$ is the output of the patchify module in DiT, obtained by a linear transformation on raw data. We plot the heat map of query and key matrices in the $5$th transformer block in Figure LABEL:fig:error-n-v(b). Plots for other transformer blocks are provided in Appendix D. As can be seen, the interaction between time embedding is dominant (bottom-right block), which aligns with our approximation theory for constructing the transformer architecture.
@@
To further understand how DiT captures the temporal dependencies, i.e., matrix $\bm{\Gamma}$ for Gaussian process data, we plot the evolution of score matrices $(\mathbf{Q}\mathbf{Y})^{\top}\mathbf{K}\mathbf{Y}$ in the attention layers at different steps in the backward diffusion process. Besides, we demonstrate the gradual change of score matrices in different attention layers. Specifically, Figure 5 presents that with progressive denoising in the backward process, the attention score matrix becomes more and more similar to the ground truth $\bm{\Gamma}$. In addition, in the first few attention layers, e.g., the first and the second layers, the dependencies are not well-structured and as predicted by our theory, these layers are still realizing some transformations on their inputs. Starting from the 3rd layer, the attention score matrices gradually exhibit the pattern of the ground truth temporal dependencies. In further subsequent layers, we observe that the learned pattern of temporal dependencies is kept. We refer readers to Figure 8 in Appendix D for a complete plot of score matrices in each attention layer.
@@ A.1 Deriving Score Function for Gaussian Process Data
$\nabla\log p_{t}(\mathbf{v}_{t})=\left(\alpha_{t}^{2}\bm{\Gamma}\otimes\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I}\right)^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu}).$
@@ A.2 Proof of Lemma 1
$\bar{\mathbf{s}}(\mathbf{v}_{t})={\left(\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)}^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu}).$
@@
¯$\bar{\mathcal{L}}_{t}(\mathbf{s})=\frac{1}{2}\mathbf{s}^{\top}{\left(\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)}\mathbf{s}-(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})^{\top}\mathbf{s},$
@@
$\left\lVert\bar{\mathbf{s}}(\mathbf{v}_{t})\right\rVert_{2}$ $={\left(\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)}^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})$
$\leq\left\lVert{\left(\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)}^{-1}\right\rVert_{2}\left\lVert\mathbf{v}_{t}-\alpha_{t}\bm{\mu}\right\rVert_{2}$
@@ B.1 Proof of Theorem 1
$\nabla\log p_{t}(\mathbf{v}_{t})=\left(\alpha_{t}^{2}(\bm{\Gamma}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu}).$ (12)
@@
$\bar{\mathbf{s}}(\mathbf{v}_{t})=\left(\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})$
@@ B.3.2 Approximate Each Transformer Subnetworks
For simplicity, for the subnetwork $f_{\rm pre}$, we assume the mean vectors $\bm{\mu}_{i}$ can be directly expressed by constant number of transformer blocks:
Assumption 2 (Construct $\bm{\mu}$).
@@
Here $L_{\mu}$ and $M_{\mu}$ are all constants. After this block, we begin to express gradient descent method using transformers. Starting at $\mathbf{s}_{i}^{(0)}=\mathbf{0}_{d}$ for $i\in[N]$, the first iteration can be written as
$\mathbf{s}_{i}^{(1)}=\mathbf{s}_{i}^{(0)}-\eta_{t}{\left(\sum_{j=1}^{N}(\alpha_{t}^{2}\bar{\bm{\Gamma}}_{ij}\bm{\Sigma})\mathbf{s}_{j}^{(0)}+\sigma_{t}^{2}\mathbf{s}_{i}^{(0)}-(\mathbf{x}_{i}-\alpha_{t}\bm{\mu}_{i})\right)}=\eta_{t}{\left(\mathbf{x}_{i}-\alpha_{t}\bm{\mu}_{i}\right)}.$
@@
where $f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})=\alpha_{t}\bm{\mu}_{i}+\bm{\epsilon}_{\mu,i}$ with $\left\lVert\bm{\epsilon}_{\mu,i}\right\rVert_{2}\leq\epsilon/\sqrt{N}$. After $f_{\rm mult}$, we could simply add one layer of transformers $\mathcal{TB}={\tt FFN}\circ{\tt Attn}$ with the attention block being trivial. We just take ${\tt FFN}(\mathbf{y})=\mathbf{y}+\mathbf{W}_{2}{\rm ReLU}(\mathbf{W}_{1}\mathbf{y})$ with
$\mathbf{W}_{1}=\begin{bmatrix}\mathbf{I}_{d}&\mathbf{0}_{d\times(4d+d_{e}+d_{t})}&-\mathbf{I}_{d}&\mathbf{0}_{d\times(3d+1)}\\ -\mathbf{I}_{d}&\mathbf{0}_{d\times(4d+d_{e}+d_{t})}&\mathbf{I}_{d}&\mathbf{0}_{d\times(3d+1)}\\ \end{bmatrix}\in\mathbb{R}^{(2d)\times D}$
@@
After this transformer block, we apply another multiplication module $f_{\rm mult,1}$ to scale the gradient by the learning rate $\eta_{t}$. We have explicitly constructed the first-step GD as $f_{\rm GD,1}=f_{\rm mult,1}\circ\mathcal{TB}\circ f_{\rm mult}$.
Size of Transformer Blocks for Approximating One GD Iteration
We summarize our construction by characterizing the size of the architecture:
  • The input to the transformer is of dimension $D\times N$ with $D=9d+d_{e}+d_{t}+1$.
  • The number of transformer blocks is bounded by $L=1+2L_{\rm mult}$. By Lemma 3, the number of layers of the multiplication module $L_{\rm mult}=\mathcal{O}(\log{\left(\left\lVert\mathbf{x}\right\rVert_{\infty}\left\lVert\mathbf{s}\right\rVert_{\infty}Nd/\epsilon\right)})$.
  • The number of attention heads is $M=1$.
  • The magnitude of the parameter in the multiplication module is at most $\mathcal{O}(\left\lVert\mathbf{x}\right\rVert_{\infty}+\left\lVert\mathbf{s}\right\rVert_{\infty})$. Since there are at most $\mathcal{O}(d)$ nonzero weights in each weight matrix, we have the norm of parameter bounded by $\mathcal{O}{\left(d(\left\lVert\mathbf{x}\right\rVert_{\infty}+\left\lVert\mathbf{s}\right\rVert_{\infty})\right)}.$
The proof is complete. ∎ The next lemma presents the construction of next $K-1$ steps of GD.
@@
$f_{{\rm GD},2}{\left(\mathbf{y}_{i}\right)}$ $=\Bigg{[}\mathbf{x}_{i}^{\top},\bm{e}_{i}^{\top},\bm{\phi}^{\top}(t),\quad\mathbf{s}_{i}^{\top}-\eta{\left(\sum_{j=1}^{N}(\alpha_{t}^{2}\bar{\bm{\Gamma}}_{ij}\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I})\mathbf{s}_{j}^{\top}-(\mathbf{x}_{i}-\alpha_{t}\bm{\mu}_{i})^{\top}\right)}+\bm{\epsilon}_{i},$
@@
$\hskip 50.00008ptM=4J,~{}~{}B=\mathcal{O}{\left(d(\left\lVert\bm{\Sigma}\right\rVert_{\infty}+r^{2}))+\left\lVert\mathbf{x}\right\rVert_{\infty}+\left\lVert\mathbf{s}\right\rVert_{\infty}\right)}.$

### v2

@@ 3.3 Bounding Correlation Truncation Error $\mathcal{E}_{2}$
Let $\mathbf{e}_{1},\dots,\mathbf{e}_{N}\in\mathbb{R}^{d_{e}}$ be $d_{e}$-dimensional time embedding of $h_{1},\dots,h_{N}$ with $\left\lVert\mathbf{e}_{i}\right\rVert_{2}=r$ for each $i$ such that there exists a positive and increasing function $f$ with $\left\lVert\mathbf{e}_{i}-\mathbf{e}_{j}\right\rVert_{2}=f(|i-j|)\geq c{\left|i-j\right|}$ for an absolute constant $c>0$. Further, the covariance function $\gamma$ satisfies
@@
Firstly, Assumption 1 says that the time embedding (a.k.a. position embedding) preserves the gap between real times. For example, a common time embedding used in sequential data modeling (Vaswani et al., 2017) is sinusoidal transformations, where $\mathbf{e}_{i}=[r\sin(2i\pi/C),r\cos(2i\pi/C)]^{\top}\in\mathbb{R}^{2}$ for a positive radius $r$ and a large constant $C>0$. We can check that $\left\lVert\mathbf{e}_{i}-\mathbf{e}_{j}\right\rVert_{2}=2r\sin({\left|i-j\right|}\pi/C)\geq 4r{\left|i-j\right|}/C$ is positive and approximately linearly increasing for a sufficiently large $C$.
Secondly, the covariance function decays exponentially fast and the speed is controlled by the exponent $\nu$ and the bandwidth $\ell$. Large $\ell$ or small $\nu$ indicates that the correlation between different time indices decays relatively slowly. Thus, the sequential data has some long-horizon dependencies. The range of $\nu$ includes the well-known quadratic-exponential (Gaussian) covariance function ($\nu=2$). Moreover, when $\nu=1$, the covariance function coincides with the correlation in a Brownian motion. Varying $\nu\in[1,2]$ can capture abundant temporal dependency patterns.
@@ 4 Approximation Theory of Score Function Using Transformers
The proof is deferred to Appendix B.1. Here we observe that the approximation error depends on $\sigma_{t}$, which matches existing results for studying score approximation using neural networks (Oko et al., 2023; Chen et al., 2023a; Tang and Yang, 2024). Yet we remark that our algorithm unrolling approach is very different from these existing works. Our results also hold for softmax activated transformer architectures, which is discussed in Appendix B.5. We remark on other interpretations.
Figure 3: Construction of score function approximation using a transformer. By rewriting the score function as the optimizer of a quadratic objective function, we use gradient descent algorithm to approximate the optimizer. We allow correlation truncation to manipulate the maximum length of temporal dependencies to model in Lemma 1. Then we construct a transformer architecture to unroll the gradient descent algorithm for score approximation in Theorem 1. Each gradient descent iteration is realized by a multiplication module followed by two transformer blocks. In the first transformer block, its attention layer calculates correlation $\bm{\Gamma}_{ij}$ utilizing time embedding. The second transformer block calculates the linear offset $-\eta_{t}\sigma_{t}^{2}\mathbf{s}-\eta_{t}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})$ in (4).
@@ 5 Sample Complexity of Diffusion Transformer
However, this loss function is not directly implementable due to $\nabla\log p_{t}$ being unknown and numerical instability when $t$ approaches zero. Therefore, we consider the following loss,
@@
Suppose Assumption 1 holds with $\ell\leq c^{\nu}$. Assume there exists a constant $C>1$ such that $C^{-1}\leq\lambda_{\min}(\bm{\Sigma})\leq\lambda_{\max}(\bm{\Sigma})\leq C$ and $\left\lVert\bm{\mu}\right\rVert_{\infty}\leq C$. We choose the transformer architecture ${\mathcal{T}}(D,L,M,B,R_{t})$ as in Theorem 1 with $\epsilon=1/n$. By setting the terminal diffusion timestep $T=\log(n)$, considering $\widehat{P}$ generated by the empirical risk minimizer $\widehat{\mathbf{s}}$, we have
@@
The proof is deferred to Appendix C. We remark that the emergence of $\widetilde{P}_{0}$ owes to the early-stopping in score estimation, which is obtained by evolving $P_{0}$ along the forward process for timestep $t_{0}$. The optimal choice on $t_{0}$ depends on the condition number $\kappa_{t_{0}}$. For fast decay covariance functions, $\kappa_{t_{0}}=\mathcal{O}(1)$ and we can choose $t_{0}=1/n$ and the generalization error is in the order of $\widetilde{\mathcal{O}}(\sqrt{\ell Nd^{3}/n})$. In this case, we have a relatively weak dependence on $N$, demonstrating the efficiency of diffusion transformers in sequential data modeling. On the other hand, when $\kappa_{t_{0}}$ is large, i.e., with the presence of long-horizon dependencies, the learning efficiency suffers from heavier dependence on $N$. We experiment with various decaying patterns in Section 6 and show the corresponding performance.
@@ 6 Numerical Results
6.1 Experiments on Gaussian Process Data
In this section, we conduct experiments on diffusion transformer for learning synthetic Gaussian process data. We test various aspects of our theory, including learning efficiency and the capture of spatial-temporal dependencies. In in Appendix D.2, we further trace the capture of spatial-temporal dependencies along with the training process and conduct additional experiments on comparing Diffusion Transformers with Unet-based Diffusion models.
@@ 6.1 Experiments on Gaussian Process Data
We consider synthetic Gaussian process data with $d=8$ and $N=128$. We stick to the setting where the covariance matrix is generated by $\bm{\Sigma}=\mathbf{A}^{\top}\mathbf{A}\in\mathbb{R}^{8\times 8}$ for a Gaussian random matrix $\mathbf{A}\in\mathbb{R}^{8\times 8}$. We set the mean vector $\bm{\mu}=\mathbf{0}$ for simplicity. The covariance function is chosen as in Assumption 1 with $\gamma(h_{i},h_{j})=\exp\left(-|h_{i}-h_{j}|^{\nu}/\ell\right).$ We set $\ell$ and $\nu$ as hyperparameters and report their influences on the learning. We generate $n\in\{1000,3200,10000,32000,100000\}$ sequences from the Gaussian process as our training data in different settings. The diffusion transformer is implemented based on the DiT (Peebles and Xie, 2023) code base. We set the number of transformer blocks to be $12$ throughout all the experiments. We modify the patchify module in DiT to cope with our Gaussian process data. Additional implementation details can be found in Appendix D.1.
@@
We dive into transformer blocks to understand how attention layer captures dependencies. Inside a transformer block, the input is a concatenation of a data vector and a corresponding time embedding written as $[\mathbf{z}_{i,t}^{\top},\mathbf{e}_{i}^{\top}]^{\top}\in\mathbb{R}^{32}$, where $\mathbf{e}_{i}\in\mathbb{R}^{16}$ is the time embedding and $\mathbf{z}_{i,t}\in\mathbb{R}^{16}$ is the output of the patchify module in DiT, obtained by a linear transformation on raw data at diffusion timestep $t$. We plot the heat map of query and key matrices in the $5$th transformer block in Figure LABEL:fig:error-n-v(b). Plots for other transformer blocks are provided in Appendix D.1. As can be seen, the interaction between time embedding (bottom-right block) is dominant, which aligns with our approximation theory for constructing the transformer architecture.
@@
To further understand how DiT captures the temporal dependencies, i.e., matrix $\bm{\Gamma}$ for Gaussian process data, we plot the evolution of score matrices $(\mathbf{Q}\mathbf{Y})^{\top}\mathbf{K}\mathbf{Y}$ in the attention layers at different steps in the backward diffusion process. Besides, we demonstrate the gradual change of score matrices in different attention layers. Specifically, Figure 5 presents that with progressive denoising in the backward process, the attention score matrix becomes more and more similar to the ground truth $\bm{\Gamma}$. In addition, in the first few attention layers, e.g., the first and the second layers, the dependencies are not well-structured and as predicted by our theory, these layers are still realizing some transformations on their inputs. Starting from the 3rd layer, the attention score matrices gradually exhibit the pattern of the ground truth temporal dependencies. In further subsequent layers, we observe that the learned pattern of temporal dependencies is kept. We refer readers to Figure 10 in Appendix D.1 for a complete plot of score matrices in each attention layer.
@@
6.2 Experiments on Semi-synthetic Video Data
To further demonstrate the capability of diffusion transformers in capturing spatial-temporal dependencies, we consider learning 2D motions of a ball. The motion is described by a sequence of gray-scale image frames of resolution $64\times 64$, which characterizes the ball that starts moving toward a random direction in a cube and bounces back when hitting a wall. Because of the bouncing-back mechanism, the dynamic of the ball goes beyond the class of Gaussian process and exemplifies more complex dependencies with abrupt changes. We train a latent diffusion model (Rombach et al., 2022), where we first generate $20000$ image frames for training a 2D Variational Autoencoder (VAE). The 2D VAE sets the latent representation in $\mathbb{R}^{2}$. Once the VAE is trained, we fix it and generate independently 10000 motion videos, each consisting of 240 image frames. The diffusion transformer is trained on the latent representations of the generated videos. As shown in Figure LABEL:fig::ball, we observe that spatial-temporal dependencies are well captured. We further collect a latent sample from the diffusion transformer and map it to the original 2D space through the pretrained decoder of the VAE, forming a $240$-frame video. We present $36$ frames in Figure 7 from the video, which are chosen uniformly out of the $240$ frames. As shown in the figure, the motion of the ball shows great time consistency and accurately captures the bouncing-back mechanism as expected.
Figure 7: Consecutive frames of a video generated by a trained diffusion transformer with a 2D VAE.
@@ Appendix A Omitted Proofs in Section 3
Before diving into detailed proofs of our results (Theorem 1 and Theorem 2), we list our assumptions for a quick reference.
$\bullet$ Data assumption. We consider Gaussian process data in $\mathbb{R}^{d}$ within interval $[0,H]$, which is defined in Section 2. The covariance function of the underlying Gaussian process verifies Assumption 1. For technical convenience, we also assume the mean function $\bm{\mu}(\cdot)$ of the Gaussian process can be efficiently represented by neural networks in Assumption 2.
$\bullet$ Transformer architecture. We denote a transformer architecture by ${\mathcal{T}}(D,L,M,B,R)$, where $D,L,M,B,R$ are hyperparameters defining the size of the transformer. In Theorem 1, we will choose these hyperparameters depending on the desired approximation accuracy. In Theorem 2, we will further choose these hyperparameters depending on the training sample size $n$.
@@ A.1 Deriving Score Function for Gaussian Process Data
$\nabla\log p_{t}(\mathbf{v}_{t})=-\left(\alpha_{t}^{2}\bm{\Gamma}\otimes\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I}\right)^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu}).$
@@ A.2 Proof of Lemma 1
$\bar{\mathbf{s}}(\mathbf{v}_{t})=-{\left(\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)}^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu}).$
@@
¯$\bar{\mathcal{L}}_{t}(\mathbf{s})=\frac{1}{2}\mathbf{s}^{\top}{\left(\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)}\mathbf{s}+(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})^{\top}\mathbf{s},$
@@
$\left\lVert\bar{\mathbf{s}}(\mathbf{v}_{t})\right\rVert_{2}$ $\leq\left\lVert{\left(\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)}^{-1}\right\rVert_{2}\left\lVert\mathbf{v}_{t}-\alpha_{t}\bm{\mu}\right\rVert_{2}$
@@ B.1 Proof of Theorem 1
$\nabla\log p_{t}(\mathbf{v}_{t})=-\left(\alpha_{t}^{2}(\bm{\Gamma}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu}).$ (12)
@@
$\bar{\mathbf{s}}(\mathbf{v}_{t})=-\left(\alpha_{t}^{2}(\bar{\bm{\Gamma}}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})$
@@ B.3.2 Approximate Each Transformer Subnetworks
For simplicity, for the subnetwork $f_{\rm pre}$, we assume that the mean vectors $\bm{\mu}_{i}$ can be directly expressed by a constant number of transformer blocks:
Assumption 2.
@@
Here $L_{\mu}$ and $M_{\mu}$ are all constants. This assumption is mild, given the approximation ability of transformers.
After applying $f_{\mu}$, we begin to implement gradient descent algorithm using transformer blocks. Starting at $\mathbf{s}_{i}^{(0)}=\mathbf{0}_{d}$ for $i\in[N]$, the first iteration can be written as
$\mathbf{s}_{i}^{(1)}=\mathbf{s}_{i}^{(0)}-\eta_{1}{\left(\sum_{j=1}^{N}(\alpha_{t}^{2}\bar{\bm{\Gamma}}_{ij}\bm{\Sigma})\mathbf{s}_{j}^{(0)}+\sigma_{t}^{2}\mathbf{s}_{i}^{(0)}+(\mathbf{x}_{i}-\alpha_{t}\bm{\mu}_{i})\right)}=-\eta_{1}{\left(\mathbf{x}_{i}-\alpha_{t}\bm{\mu}_{i}\right)}.$
@@
where $f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})=\alpha_{t}\bm{\mu}_{i}+\bm{\epsilon}_{\mu,i}$ with $\left\lVert\bm{\epsilon}_{\mu,i}\right\rVert_{2}\leq\epsilon/\sqrt{N}$. After obtaining $f_{\rm mult}$, we use one transformer block $\mathcal{TB}={\tt FFN}\circ{\tt Attn}$ with the attention block being trivial, i.e., all weight parameters being zero so that ${\tt Attn}(\mathbf{y})=\mathbf{y}$. We then choose ${\tt FFN}(\mathbf{y})=\mathbf{y}+\mathbf{W}_{2}{\rm ReLU}(\mathbf{W}_{1}\mathbf{y})$ with
$\mathbf{W}_{1}=\begin{bmatrix}-\mathbf{I}_{d}&\mathbf{0}_{d\times(4d+d_{e}+d_{t})}&\mathbf{I}_{d}&\mathbf{0}_{d\times(3d+1)}\\ \mathbf{I}_{d}&\mathbf{0}_{d\times(4d+d_{e}+d_{t})}&-\mathbf{I}_{d}&\mathbf{0}_{d\times(3d+1)}\\ \end{bmatrix}\in\mathbb{R}^{(2d)\times D}$
@@
We further apply another multiplication module $f_{\rm mult,1}$ to rescale the gradient computed by the transformer block by the learning rate $\eta_{1}$. We can check that the first-step GD is represented by $f_{\rm GD,1}=f_{\rm mult,1}\circ\mathcal{TB}\circ f_{\rm mult}$.
Size of Transformer Blocks for Approximating The First GD Iteration
We summarize in the following table the resulting network size of transformer blocks for implementing the first GD iteration when initialized with $\mathbf{s}_{i}=\bm{0}$.
Table 1: Transformer size for approximating the first GD iteration
Input dimension $D\times N$ with $D=9d+d_{e}+d_{t}+1$
# of blocks $L$ $1+2L_{\rm mult}$ with $L_{\rm mult}=\mathcal{O}(\log{\left(\left\lVert\mathbf{x}\right\rVert_{\infty}\left\lVert\mathbf{s}\right\rVert_{\infty}Nd/\epsilon\right)})$
# of attention heads $M$ 1
Parameter bound $B$ $\mathcal{O}{\left(d(\left\lVert\mathbf{x}\right\rVert_{\infty}+\left\lVert\mathbf{s}\right\rVert_{\infty})\right)}$
Here the parameter bound $B$ is obtained by noting that the magnitude of weight parameters in the multiplication module is at most $\mathcal{O}(\left\lVert\mathbf{x}\right\rVert_{\infty}+\left\lVert\mathbf{s}\right\rVert_{\infty})$ and there are at most $\mathcal{O}(d)$ nonzero weight parameters in each weight matrix. The proof is complete. ∎ The next lemma presents the construction of next $K-1$ steps of GD.
@@
$f_{{\rm GD},2}{\left(\mathbf{y}_{i}\right)}$ $=\Bigg{[}\mathbf{x}_{i}^{\top},\bm{e}_{i}^{\top},\bm{\phi}^{\top}(t),\quad\mathbf{s}_{i}^{\top}-\eta_{t}{\left(\sum_{j=1}^{N}(\alpha_{t}^{2}\bar{\bm{\Gamma}}_{ij}\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I})\mathbf{s}_{j}^{\top}+(\mathbf{x}_{i}-\alpha_{t}\bm{\mu}_{i})^{\top}\right)}+\bm{\epsilon}_{i}^{\top},$
@@
$D=9d+d_{e}+d_{t}+1,\quad L=\mathcal{O}(\log{\left(\left\lVert\mathbf{x}\right\rVert_{\infty}\left\lVert\mathbf{s}\right\rVert_{\infty}Nd/\epsilon\right)}),$
$M=4J,\quad B=\mathcal{O}{\left(d(\left\lVert\bm{\Sigma}\right\rVert_{\infty}+r^{2}))+\left\lVert\mathbf{x}\right\rVert_{\infty}+\left\lVert\mathbf{s}\right\rVert_{\infty}\right)}.$